# Joint Problems in Learning Multiple Dynamical Systems

### v3

@@ 2 Background
2.1 Linear Dynamic Systems (LDS) and System Identification
@@
2.2 Clustering with LDS Assumptions
@@
2.3 Mixed-Integer Programming (MIP)
@@
While MINLP problems may seem too general a model for our joint problem, notice that the NP-hardness and inapproximability results discussed in Section 2.1 suggest that this may be the appropriate framework. For bounded variables, standard branch-and-bound-and-cut algorithms run in finite time. Both in theory – albeit under restrictive assumptions, such as in [17] – and in practice, the expected runtime is often polynomial. In the formulation of the next section, $f,g_{i}$ are trilinear, and various monomial envelopes have been considered and implemented in global optimization solvers such as BARON [48], SCIP [6], and Gurobi  [^1] https://www.gurobi.com/ [22]. In our approach, we consider a mixed-integer programming formulation of a piecewise polyhedral relaxation of a multilinear term using its convex-hull representation.
2.4 Non-Commutative Polynomial Optimization (NCPOP)
@@
[…]
@@
Under the Archimedean assumption, such that the tuple of operators $X$ are bounded, one can utilize the Sums of Squares theorem of [24] and [39] to derive semidefinite programming (SDP) relaxations of the Navascules-Pironio-Acin (NPA) hierarchy [45]. There are also variants [55, 54, 53] that exploit various forms of sparsity.
3 Problem Formulation
@@
3.1 Least-Squares Objective Function
@@
In the first part of the optimization criterion (35), we have a sum of squares of the difference between trajectory estimate $f_{t}^{l_{i}}$ and observations of the trajectories assigned to cluster $C_{c}$ for $c=0,1$. The second part of the optimization criterion (9) provides a form of regularization.
3.2 Feasible Set in State Space
@@
The first two equations in the set of constraints, (10) and (11), encode that the system is an LDS with system matrices $\mathbf{F}$ and $\mathbf{G}$. For brevity of the notation, both equations are indexed by the cluster index $c$, which can be rewritten as twice as many equations, one with $f^{0}_{t},\mathbf{F}_{0},\mathbf{G}_{0},\upsilon^{0}_{t},\omega^{0}_{t}$, and $\phi_{t}^{0}$, the second one with $f^{1}_{t},\mathbf{F}_{1},\mathbf{G}_{1},\upsilon^{1}_{t},\omega^{1}_{t}$, and $\phi_{t}^{1}$. The third equation (12) encodes that the indicator function is 00 or $1$ for each trajectory.
or even a weighted combination of the redundant constraints in the spirit of Gomory. While these strengthen the relaxations, the higher-degree polynomials involved come at a considerable cost. Still, even when the dimension of the hidden state n is unknown, we can solve the corresponding operator-valued problem:
@@ 3.3 Variants and Guarantees
4 EM Heuristic
In addition to tackling the optimization problem in Section 3.1, we provide an alternative solver using the Expectation-Maximization (EM) heuristic [16]. The main idea of the algorithm is to avoid the direct optimization of the criterion in (35). Instead, the indicator function is treated separately. In the expectation step, the parameters of the LDSs are fixed, and the assignment of the trajectories to the clusters (i.e., the indicator function) is calculated. Then, in the maximization step, the criterion is optimized, and the LDS parameters are calculated with the indicator function fixed. See Algorithm 1 for the pseudocode. First, the algorithm randomly partitions the trajectories into the clusters. Then, for each cluster, the parameters of the LDSs are found, and with the parameters known, the optimization criterion is recalculated, and each trajectory is put to the cluster, which lowers the error in (35).
@@ 4 EM Heuristic
As EM-algorithm applied to the mixture of Gaussians is, in many scenarios, computationally inefficient, we propose to use heuristic 1, which can be seen as a parallel to the Lloyd’s algorithm [37] for $k$-means problem. See the Supplementary materials for a formal connection to the $k$-means problem.
function EM-clustering($N$ trajectories $\mathbf{X}^{i}\in\mathbb{R}^{m\times T}$, $K$)
@@ 5.1 Methods and Solvers
The discussion of runtime As a subroutine of the EM Heuristic, the LDS identification in equation (35) is solved via Bonmin, with runtime presented in the center subplot of Figure 3. When the dimensions of system matrices are not assumed, the LDS identification becomes an NCPOP, and the runtime increases exponentially if NPA hierarchy [45] is used to find the global optimal solutions, but stays relatively modest if sparity exploit variants [55, 54, 53] are used.
@@
  • Dynamic Time Warping (DTW) is used for measuring similarity between given time series [49, 21]. We utilize K-means on DTW distance with tslearn [JMLR:v21:20-091] package.
@@
5.2 Experiments on Synthetic Data
@@ Results
5.3 Experiments on Real-world Data
@@
Table 1: F1 score of EM Heuristic using NCPOP compared with the method requiring specific dimension $n$ of hidden state with window size $T$ in $\{10,20,30\}$. $n$ is chosen from $\{2,3,4\}$.
Finally, when the dimension $n$ of the hidden state is not assumed, the subproblem of the EM heuristic becomes an NCPOP. For the implementation with such an assumption, we construct NCPOP using ncpol2sdpa 1.12.2444https://ncpol2sdpa.readthedocs.io/en/stable/ [59]. Subsequently, the relaxation problem is solved by Mosek 10.1555https://www.mosek.com/ [41]. Noted that the execution time of NCPOP escalates rapidly as the trajectory length $T$ grows, we test NCPOP and compare its performance with the aforementioned EM Heuristic method with $T\in\{10,20,30\}$. For comparison, we use pyomo [^6] https://www.pyomo.org/ [11] to construct the model and solve the problem with Bonmin [^7] https://www.coin-or.org/Bonmin/ [8], as above, with the dimension $n$ from $\{2,3,4\}$. The overall performance is illustrated in Table 1. NCPOP demonstrates the best performance in terms of the $F_{1}$ score when $T=10$ and $30$. However, as shown in the right subplots of Figure 3, the runtime of NCPOP grows significantly as the length $T$ of the trajectory increases.
6 Conclusions and Further Work
We have studied problems in clustering time series, where given a set of trajectories and a number of parts, we jointly partition the set of trajectories and estimate a linear dynamical system (LDS) model for each part, so as to minimize the maximum error across all the models. As discussed in Section 3.3, a number of variants of the joint problem remain to be investigated. The computational aspects of the operator-valued problem [66] that consider the dimension of the hidden state to be unknown seem particularly interesting.
Appendix A Table of Notation Used in the Paper
Table 2: A Table of Notation.
@@ B.1 Preliminaries
$\mathcal{N}\left(\begin{bmatrix}\mu_{x}\\ \mu_{y}\end{bmatrix},\begin{bmatrix}\Sigma_{x}&0\\ 0&\Sigma_{y}\end{bmatrix}\right)$
@@
$\begin{pmatrix}x\\ \mathbf{A}x+y\end{pmatrix}\sim\mathcal{N}\left(\begin{pmatrix}\mu_{x}\\ \mathbf{A}x\end{pmatrix},\begin{bmatrix}\Sigma_{x}&\Sigma_{x}\mathbf{A}^{\prime}\\ \mathbf{A}\Sigma_{x}&\mathbf{A}\Sigma_{x}\mathbf{A}^{\prime}+\Sigma_{y}\end{bmatrix}\right).$ (13)
@@
$\begin{pmatrix}x\\ y\end{pmatrix}\sim\mathcal{N}\left(\begin{pmatrix}\mu_{x}\\ 0\end{pmatrix},\begin{bmatrix}\Sigma_{x}&0\\ 0&\Sigma_{y}\end{bmatrix}\right).$ (14)
@@ B.2 Analysis of the EM-algorithm
B.3 Implications of the Normally Distributed Observations
@@
B.4 Practical Applicability of the Gaussian Mixture-Based EM-algorithm
@@
function EM-clustering($N$ trajectories $x_{t}^{i}\in\mathbb{R}^{m\times T}$, $K$)
@@ B.5 Connection to $k$-means
Appendix C Clustering Performance Metrics
@@
where precision $=\frac{TP}{TP+FP}$ and recall $=\frac{TP}{TP+FN}$ and $TP$, $FP$, and $FN$ are the numbers of true positives, false positives, and false negatives, respectively. We calculate the $F_{1}$ score twice for each class, once with one class labeled as positive and once with the other class labeled as positive, and then we select the higher score for each class. This approach is used because there is no predefined positive and negative labels.

### v4

# Mengjia Niu Imperial College London South Kensington Campus London SW7 2AZ, UK

@@ II Background
II-A Linear Dynamic Systems (LDS) and System Identification
@@
II-B Clustering with LDS Assumptions
@@
II-C Mixed-Integer Programming (MIP)
@@
While MINLP problems may seem too general a model for our joint problem, notice that the NP-hardness and inapproximability results discussed in Section II-A suggest that this may be the appropriate framework. For bounded variables, standard branch-and-bound-and-cut algorithms run in finite time. Both in theory – albeit under restrictive assumptions, such as in [18] – and in practice, the expected runtime is often polynomial. In the formulation of the next section, $f,g_{i}$ are trilinear, and various monomial envelopes have been considered and implemented in global optimization solvers such as BARON [53], SCIP [7], and Gurobi. In our approach, we consider a mixed-integer programming formulation of a piecewise polyhedral relaxation of a multilinear term using its convex-hull representation.
II-D Non-Commutative Polynomial Optimization (NCPOP)
@@
Let $X=(X_{1},\ldots,X_{n})$ be a tuple of bounded operators on a Hilbert space $\mathcal{H}$. Let $X,X^{{\dagger}}]$ denote these $2n$ operators, with the $\dagger$}-algebra being conjugate transpose. Let monomials $\omega,\mu$ be products of powers of variables from $X,X^{{\dagger}}]$. The degree of a monomial, denoted by $|\omega|$, refers to the sum of the exponents of all operators in the monomial $\omega$, e.g., $|X^{3}_{n}X_{n}^{{\dagger}}|=4$. Let $p$ and $q_{i}$, $i=1,\dots,m$ be polynomials in these $2n$ variables. Let $\deg(p)$ denote the polynomial degree of $p$. In the following, we will view these $2n$ variables as the new tuple $X$. Using the set of monomials generated from the tuple $X$, polynomials $p$ and $q_{i}$, $i=1,\dots,m$ can be rewritten as linear combinations of monomials:
@@
Under the Archimedean assumption, such that the tuple of operators $X$ are bounded, one can utilize the Sums of Squares theorem of [26] and [42] to derive semidefinite programming (SDP) relaxations of the Navascules-Pironio-Acin (NPA) hierarchy [46, 50]. There are also variants [62, 61, 60] that exploit various forms of sparsity.
III Problem Formulation
@@
III-A Least-Squares Objective Function
@@
In the first part of the optimization criterion (35), we have a sum of squares of the difference between trajectory estimate $f_{t}^{l_{i}}$ and observations of the trajectories assigned to cluster $C_{c}$ for $c=0,1$. The second part of the optimization criterion (9) provides a form of regularization with parameter=$270$.
III-B Feasible Set in State Space
@@
The first two equations in the set of constraints, (10) and (11), encode that the system is an LDS with system matrices $\mathbf{F}$ and $\mathbf{G}$. For brevity of the notation, both equations are indexed by the cluster index $c$, which can be rewritten as twice as many equations, one with $f^{0}_{t},\mathbf{F}_{0},\mathbf{G}_{0},\upsilon^{0}_{t},\omega^{0}_{t}$, and $\phi_{t}^{0}$, the second one with $f^{1}_{t},\mathbf{F}_{1},\mathbf{G}_{1},\upsilon^{1}_{t},\omega^{1}_{t}$, and $\phi_{t}^{1}$. The third equation (12) encodes that the indicator function is 0 or $1$ for each trajectory.
A weighted combination of the redundant constraints in the spirit of Gomory. While these strengthen the relaxations, the higher-degree polynomials involved come at a considerable cost. Still, even when the dimension of the hidden state n is unknown, we can solve the corresponding operator-valued problem:
@@ III-C Variants and Guarantees
IV EM Heuristic
In addition to tackling the optimization problem in Section III-A, we provide an alternative solver using the Expectation-Maximization (EM) heuristic [17]. The main idea of the algorithm is to avoid the direct optimization of the criterion in (35). Instead, the indicator function is treated separately. In the expectation step, the parameters of the LDSs are fixed, and the assignment of the trajectories to the clusters (i.e., the indicator function) is calculated. Then, in the maximization step, the criterion is optimized, and the LDS parameters are calculated with the indicator function fixed. See Algorithm 2 for the pseudocode. First, the algorithm randomly partitions the trajectories into the clusters. Then, for each cluster, the parameters of the LDSs are found, and with the parameters known, the optimization criterion is recalculated, and each trajectory is put to the cluster, which lowers the error in (35).
@@ IV EM Heuristic
As EM-algorithm applied to the mixture of Gaussians is, in many scenarios, computationally inefficient, we propose to use heuristic 2, which can be seen as a parallel to the Lloyd’s algorithm [40] for $k$-means problem. See the Supplementary materials for a formal connection to the $k$-means problem.
function EM-clustering(NN trajectories $\mathbf{X}^{i}\in\mathbb{R}^{m\times T}$, $K$)
@@ V-A Methods and Solvers
The discussion of runtime As a subroutine of the EM Heuristic, the LDS identification in equation (35) is solved via Bonmin, with runtime presented in the center subplot of Figure 3. When the dimensions of system matrices are not assumed, the LDS identification becomes an NCPOP, and the runtime increases exponentially if NPA hierarchy [46, 50] is used to find the global optimal solutions, but stays relatively modest if sparity exploit variants [62, 61, 60] are used.
@@
  • Dynamic Time Warping (DTW) is used for measuring similarity between given time series [54, 23]. We utilize K-means on DTW distance with tslearn [56] package.
@@
V-B Experiments on Synthetic Data
@@ Results
V-C Experiments on Real-world Data
@@
TABLE I: F1 score of EM Heuristic using NCPOP compared with the method requiring specific dimension $n$ of hidden state with window size $T$ in $\{10,20,30\}$. $n$ is chosen from $\{2,3,4\}$.
Finally, when the dimension $n$ of the hidden state is not assumed, the subproblem of the EM heuristic becomes an NCPOP. For the implementation with such an assumption, we construct NCPOP using ncpol2sdpa 1.12.2333https://ncpol2sdpa.readthedocs.io/en/stable/ [66]. Subsequently, the relaxation problem is solved by Mosek 10.1444https://www.mosek.com/ [45]. Noted that the execution time of NCPOP escalates rapidly as the trajectory length $T$ grows, we test NCPOP and compare its performance with the aforementioned EM Heuristic method with $T\in\{10,20,30\}$. For comparison, we use pyomo [^5] https://www.pyomo.org/ [12] to construct the model and solve the problem with Bonmin [^6] https://www.coin-or.org/Bonmin/ [9], as above, with the dimension $n$ from $\{2,3,4\}$. The overall performance is illustrated in Table I. NCPOP demonstrates the best performance in terms of the $F_{1}$ score when $T=10$ and $30$. However, as shown in the right subplots of Figure 3, the runtime of NCPOP grows significantly as the length $T$ of the trajectory increases.
VI Conclusions and Further Work
We have studied problems in clustering time series, where given a set of trajectories and a number of parts, we jointly partition the set of trajectories and estimate a linear dynamical system (LDS) model for each part, so as to minimize the maximum error across all the models. As discussed in Section III-C, a number of variants of the joint problem remain to be investigated. The computational aspects of the operator-valued problem [73] that consider the dimension of the hidden state to be unknown seem particularly interesting.
Supplementary Materials
I Properties of Presented Method
Considering the above contributions with previous work produced by [37, 29, 43], a pioneering method can be created that holds following properties:
  • Mimic temporal dynamics in the data and produces the related, interpretable features, unlike methods relying on deep learning.
  • Handle multiple time lags within the time series without losing its periodicity, unlike methods based on Fourier coefficients [31].
  • Not necessary to make any linear-algebraic assumptions, unlike subspace methods [2, 19] that assume the existence of a subspace, which might not necessarily be true. The assumption that the time series were generated by an LDS does not cause any loss of generality.
  • Avoid of making any assumptions on the dimension of the hidden state, unlike previous methods utilizing LDS [37, 38, 5, 43].
  • Improves upon previously proposed methods based on Fast Fourier Transform (FFT) [31] and Dynamic Time Warping (DTW) [54, 23, 55] in practical experiments of Section V.
II Table of Notation Used in the Paper
TABLE II: A Table of Notation.
@@ III-A Preliminaries
[…]
@@
[…]
@@
[…]
@@ III-B Analysis of the EM-algorithm
III-C Implications of the Normally Distributed Observations
@@
III-D Practical Applicability of the Gaussian Mixture-Based EM-algorithm
@@
function EM-clustering(NN trajectories $x_{t}^{i}\in\mathbb{R}^{m\times T}$, $K$)
@@ III-E Connection to $k$-means
IV Clustering Performance Metrics
@@
where precision $\frac{TP}{TP+FP}$ and recall $\frac{TP}{TP+FN}$ and $TP$, FPFP, and $FN$ are the numbers of true positives, false positives, and false negatives, respectively. We calculate the $F_{1}$ score twice for each class, once with one class labeled as positive and once with the other class labeled as positive, and then we select the higher score for each class. This approach is used because there is no predefined positive and negative labels.